Suppose a monopolisti has a demand curve that can be expressed as P -19 minus Q. Turn on the police marginal revenue curve can be expressed as MR-90-2Q.
The monopolist has constant marginal cost and average total cost of $10. Find the deadweight loss of a profit maximizing monopolist
Answer #1

Please note that, you have given the demand curve is P = 19-Q, at this demand price at profit maximising comes in negative which is not possible. Therefore I have taken the demand curve P = 90-Q. If you have any doubts, kindly ask in comment.

solution- Given; P= too MRE 90-0 go- 20 mc = ATC= 10 we knows that for profit maximization; MR = mc go- 2a = 10 zs = 80 0 = 4